OpenStudy (anonymous):

plz help! list the discontinuities for the function f(x)=cot(5x/3) choices: no discontinuities n(pi/2), where n is an integer n(pi/3),where n is an integer n(3pi/5) where n is an integer

OpenStudy (chillout):

\[\cot(5x/3)=\frac{ \cos(5x/3) }{ \sin(5x/3) }\]find the values in which sin(5x/3)=0.

OpenStudy (anonymous):

C?

OpenStudy (chillout):

Yep.

OpenStudy (michele_laino):

hint: this is the graph of cot(x):